# LoSA: Long-Short-range Adapter for Scaling End-to-End Temporal Action Localization

### v1

@@ 5 Conclusion
We introduce LoSA, the first memory-and-parameter-efficient backbone adapter designed specifically for TAL. LoSA comprises a novel design of Long-range and Short-range Temporal Adapters that are attached to the intermediate layers to adapt them towards improving TAL, and run parallel to the video backbone to reduce memory footprint. Finally, Long-Short-range Fusion module takes the output from these adapters to fuse and give TAL-enhanced features. This allows LoSA to scale end-to-end backbone adaptation to billion-parameter-plus backbones like VideoMAEv2 (ViT-g) and leverage them beyond head-only training to significantly outperform all existing TAL methods. Our work is the first to go beyond traditional techniques, including full model adaptation and head-only transfer learning, in addressing the challenging problem of adapting video backbones for end-to-end TAL, and proves effective in leveraging large foundation models for performing action localization in untrimmed videos.
@@ 6 Acknowledgment
S2 Additional Analysis
In Fig S3, we conduct sensitivity analysis for THUMOS-14 using VideoMAEv2-(VIT-g). This chart takes into consideration coverage, length and number of instances. Coverage presents the relative length of actions compared to the whole video and is further categorized into five bins: Extra Small (XS: (0, 0.02]), Small (S: (0.02, 0.04]), Medium (M: (0.04, 0.06]), Large (L:(0.06, 0.08]), and Extra Large (XL: (0.08, 1.0]). Length presents the actual length of actions and is further divided into five groups: Extra Small (XS: (0, 3]), Small (S: (3, 6]), Medium (M: (6, 12]), Long (L: (12, 18]), and Extra Long (XL: > 18). Finally, number of instances of same class in a video are divided into four parts: Extra Small (XS: 1); Small (S: [2, 40]); Medium (M: [40, 80]); Large (L: > 80). In Fig S3, we show comparison between the baseline, LoSA w/o Long-Short-range Adapter (the top row) and our method LoSA (the bottom row). We can observe that the performance for both XS and XL improves significantly with our method LoSA compared to the baseline, LoSA w/o Long-Short-range Adapter. We believe this is because by having the untrimmed temporal learning at different temporal ranges in the intermediate layers via the long-range and short-range temporal adapters, LoSA is able to capture complex scene details required for long action instances as well as fine-grained information required for shorter duration actions respectively.
In Fig S4, we conduct a False Positive (FP) analysis at tIoU=0.5 for THUMOS-14 using VideoMAEv2 (VIT-g). Similar to Fig S3, we show comparison between the baseline, LoSA w/o Long-Short-range Adapter (Fig S4a) and our method LoSA (Fig S4b). We can see a drop in the wrong label prediction error with LoSA compared to LoSA w/o Long-Short-range Adapter. This shows the significance of incorporating untrimmed temporal video understanding while adapting the intermediate layers for TAL. The chart shows FP error breakdown for top-10 ground truth (GT) predictions. For more details regarding the chart, we refer the readers to [2].
Refer to captionRefer to caption
Figure S3: Sensitivity analysis on THUMOS-14 using [2]. mAPN denotes normalized mAP at tIoU=0.5 with N average ground truth segments per class. Top: LoSA w/o Long-Short-Adapter. Bottom: LoSA (ours). Performance for both XS and XL improves significantly with our method LoSA (bottom) compared to the baseline, LoSA w/o Long-Short-range Adapter (top).
@@
S3 Limitations, Negative Impact, and Future Work

### v2

@@ 5 Conclusion
We introduce LoSA, the first memory-and-parameter-efficient backbone adapter designed specifically for TAL. LoSA comprises a novel design of Long-range and Short-range Temporal Adapters that are attached to the intermediate layers to adapt them towards improving TAL, and run parallel to the video backbone to reduce memory footprint. Finally, Long-Short-range Gated Fusion module takes the output from these adapters to fuse and give TAL-enhanced features. This allows LoSA to scale end-to-end backbone adaptation to $>$1 B parameter backbones like VideoMAEv2 (ViT-g) and leverage them beyond head-only training to significantly outperform all existing TAL methods. Our work is the first to go beyond traditional techniques, including full model adaptation and head-only transfer learning, in addressing the challenging problem of adapting video backbones for end-to-end TAL, and proves effective in leveraging large foundation models for TAL in untrimmed videos.
@@ 6 Acknowledgment
S3 Additional Analysis
In Fig S3, we conduct a False Positive (FP) analysis at tIoU=0.5 for THUMOS-14 using VideoMAEv2 (VIT-g). We show comparison between the baseline, LoSA w/o Long-Short-range Adapter (Fig S3a) and our method LoSA (Fig S3b). We can see a drop in the wrong label prediction error with LoSA compared to LoSA w/o Long-Short-range Adapter. This shows the significance of incorporating untrimmed temporal video understanding while adapting the intermediate layers for TAL. The chart shows FP error breakdown for top-10 ground truth (GT) predictions. For more details regarding the chart, we refer the readers to [2].
@@
S4 Limitations, Negative Impact, and Future Work